let B={1-t,t-t2,2-2t+t2} be an ordered basis for p2

let p(t)=3+t-6t2

  1. find [p(t)]B
1
Expert's answer
2022-04-26T05:34:25-0400

Denote: "p_1=1-t", "p_2=t-t^2", "p_3=2-2t+t^2". The aim is to find real numbers "\alpha_1,\alpha_2,\alpha_3" satisfying: "p(t)=\alpha_1p_1+\alpha_2p_2+\alpha_3p_3". We get the following equations for "\alpha_1", "\alpha_2" and "\alpha_3" by considering coefficients near "t^2", "t", "1": "-6=\alpha_3-\alpha_2", "1=-\alpha_1+\alpha_2-2\alpha_3" and "3=\alpha_1+2\alpha_3". From the first and third equation we get: "\alpha_2=\alpha_3+6", "\alpha_1=3-2\alpha_3". By substituting it in the second equation we get: "1=-(3-2\alpha_3)+(\alpha_3+6)-2\alpha_3". Thus, "\alpha_3=-2", "\alpha_2=4", "\alpha_1=7".

Answer: "p(t)=7(1-t)+4(t-t^2)-2(2-2t+t^2)".